Internal Analysis of Marriott- calculations of relevant profit ratios, activity ratios, leverage ratios, liquidity ratios, and market ratios and any crucial financial information.  
Answer #1

I will list the required ratios below based on the latest Audited financial results available i.e. Year ending Dec31, 2017:

All figures are in $million.

Profit Ratios:

1) Return on Assets % = EBIT x (1 - 0.375) / Total assets + Total assets (Previous year) /2

Return on Assets % = 2518 x (1-0.375) / 23948 + 24140 /2

Return on Assets % = 6.5%

2) Return on Equity % = EBITDA / Total equity + Total equity (Previous year) /2

Return on Equity % = 1372 / 3731 + (5753 / 2) = 30.2%

Activity ratios:

1) Total asset turnover = Total revenues / Total assets + Total assets (Previous year) /2

Total asset turnover = 5129 / 23948 + (24140/2) = 0.21

Leverage ratios:

1) Debt / equity ratio = Total debt / Total equity = 8238 / 3731 = 220.8%

2) Debt / Capital = Total debt / Total capital = 8238 / 11969 = 68.8%

Liquidity ratios:

1) Current ratio = Current assets / Current liabilities = 2747 / 6010 =0.46

2) Quick ratio = Cash and short term investments + Accounts receivables / total current liabilities

Quick ratio = 383 + 1984 / 6010 = 0.39

Market ratios:

1) Price to BV ratio = Stock price / Book value per share = 15.71